# 2]organization=Shanghai Blockchain Engineering Research Center, city=Shanghai, postcode=200433, country=China

### v1

@@ 1 Function StealthAttack($H^{\text{target}}$, $D^{\text{priv}}$):
3      $g_{\text{target}}$=TrainInverseNet2($H^{\text{target}}$ , $D^{\text{priv}}$)
4      $\hat{x}_{\text{target}}$ = Inverse($g_{\text{target}}$, $H^{\text{target}}$)
@@
8Function TrainInverseNet2($H^{\text{target}}$, $D^{\text{priv}}$):
@@ 10      $g_{\text{target}}^{(0)}=$ Init()
13            Obtain leaked samples $x_{1},x_{2}...x_{k}$ from $D^{\text{priv}}$
14            Prepare $H_{i}$ data corresponding to $x_{i}$
15            $L(g_{\text{target}})=\frac{1}{m}\sum_{i=1}^{m}\left\|g_{\text{target}}(H_{i})-x_{i}\right\|^{2}$
16            $\theta_{g}^{(n+1)}=\theta_{g}^{(n)}-\epsilon\frac{\partial L(g_{\text{target}}^{(n)})}{\partial\theta_{g}^{(n)}}$
@@ 17            $n+=1$
19      return $g_{\text{target}}^{(\text{NIters})}$
@@ 6.1.1 Datasets
In our experiments, we employed four public datasets: Bank marketing analysis Moro et al. (2014) [^1] https://archive.ics.uci.edu/dataset/222/bank+marketing (Bank), Adult income Becker and Kohavi (1996) [^2] https://archive.ics.uci.edu/dataset/2/adult (Adult), Default of credit card clients Yeh and Lien (2009) [^3] https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients (Credit) and CIFAR10 Krizhevsky et al. (2009) [^4] https://www.cs.toronto.edu/~kriz/cifar.html, to evaluate our methods. These datasets range from banking marketing analysis to image recognition, each with its unique features and challenges. For data preparation, continuous columns in tabular datasets were scaled, and discrete columns were one-hot encoded. We have summarized the evaluated datasets in Table 2.
Dataset Bank Adult Credit CIFAR10
@@
The Bank dataset, derived from a Portuguese banking institution’s direct marketing campaigns, encompasses 41188 samples with 20 features and two classes to determine the likelihood of clients subscribing to a term deposit, featuring inputs like age, job, marital status, and education. The Adult dataset, also known as the "Census Income" dataset, includes 32561 instances with 14 features such as age, workclass, and education, and it aims to predict if an individual’s income surpasses \$50,000 per annum using census data. The Credit dataset, sourced from Taiwan, comprises 30,000 instances with 23 features, including credit amount, gender, education, and marital status, and aims to predict the probability of default payments in credit card clients using various data mining methods. The CIFAR10 dataset is a widely-used public dataset for computer vision research, containing 60,000 color images with a resolution of 32x32 pixels, divided into ten categories with 6,000 images in each category. Additionally, before training the models, we scaled the continuous columns in the tabular datasets to the range of [-1, 1], while the discrete columns were encoded using one-hot encoding.
@@ 6.1.2 Models
The InverNet for all bottom models is consistent with the architecture of the respective bottom model. For tabular data, the inverse net utilizes a three-layer fully connected neural network, while for image data, the model employs two transposed convolutional layers and one pooling layer.
To ascertain the effectiveness of our proposed method, on the tabular data set, we conducted comparative evaluations against two types of Vertical Federated Learning (VFL) data reconstruction attacks. Firstly, we compared our approach to a Generative Model Reconstruction Attack under a white-box scenario, denoted as GRN Luo et al. (2021). Secondly, we assessed against a black-box scenario where a shadow model is trained prior to launching a white-box attack, referred to as GIA Jiang et al. (2022). Lastly, we included a random guessing baseline method to benchmark the inherent performance of random speculation. On the image dataset, we compared our method with split learning data reconstruction attack approaches He et al. (2019), namely Black-box attack and Query-free attack methods.
In pursuit of a fair comparison with other methodologies, our model’s architecture mirrors that of GRN Luo et al. (2021) in terms of the number of parameters, featuring three hidden layers with neuron counts of 600, 300, and 100. We designed our individual bottom models with 300, 100, and 100 neurons in their hidden layers. We also took care to ensure that the GIA method Jiang et al. (2022) we implemented utilized an equivalent number of model parameters for an accurate comparison. Moreover, despite the original GIA method not incorporating nonlinear activation functions, we introduced them into the target party’s model to ensure the fairness of the comparison. For the reconstruction method of image data, we use the same model architecture. However, unlike the split learning model, the VFL has two separate bottom models that are trained independently.
@@ 6.2 Evaluation Metrics
For image data, we adopt two widely recognized metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) Wang et al. (2004). PSNR quantifies image errors by calculating the mean squared error between origin and attack images, with higher values indicating lower quality degradation. SSIM evaluates image quality based on structural information, brightness, and contrast, ranging from 0 to 1, with 1 indicating perfect similarity.
In evaluating tabular data, previous studies Luo et al. (2021, 2022) have used training loss or distance measures to assess reconstruction accuracy. However, these methods may not align with real attack scenarios, which focus on whether reconstructed categories match the actual ones. To address these issues, we adopted the metrics proposed in Vero et al. (2022). Considering the characteristics of tabular data, we separate the treatment of categorical and continuous features. For vector $x$ and its reconstruction vector $\hat{x}$, the accuracy metric is defined as follows:
$\text{accuracy}(x,\hat{x}):=$ $\frac{1}{M+L}\bigg{(}\sum_{i=1}^{M}\mathbb{I}\{x^{D}_{i}=\hat{x}^{D}_{i}\}$
$+\sum_{i=1}^{L}\mathbb{I}\{\hat{x}^{C}_{i}\in[x^{C}_{i}-\varepsilon,x^{C}_{i}+\varepsilon]\}\bigg{)},$ (6)
@@ 6.3 Performance Evaluation and Comparison
Dataset Evaluation UIFV-QA UIFV-DPA UIFV-IQA UIFV-SA GRN GIA Random
Bank Accuracy $98.0\pm 0.1$ $95.7\pm 0.4$ $66.2\pm 1.4$ $90.1\pm 0.1$ $30.2\pm 6.7$ $55.9\pm 1.8$ $21.0\pm 0.0$
Discrete Acc $99.5\pm 0.0$ $99.0\pm 0.1$ $87.8\pm 2.1$ $94.9\pm 0.3$ $29.4\pm 8.0$ $82.6\pm 4.5$ $26.9\pm 0.1$
Continuous Acc $96.4\pm 0.2$ $92.4\pm 0.8$ $44.6\pm 1.2$ $85.2\pm 0.3$ $30.5\pm 10.5$ $29.3\pm 3.3$ $15.1\pm 0.1$
Adult Accuracy $98.5\pm 0.0$ $80.8\pm 0.9$ $94.8\pm 0.5$ $72.8\pm 3.4$ $41.1\pm 8.5$ $18.9\pm 8.3$ $11.1\pm 0.1$
Discrete Acc $98.4\pm 0.0$ $81.1\pm 0.9$ $98.1\pm 0.6$ $71.4\pm 2.4$ $41.9\pm 6.4$ $23.4\pm 10.1$ $9.0\pm 0.1$
Continuous Acc $98.8\pm 0.1$ $79.8\pm 1.2$ $81.7\pm 0.7$ $78.3\pm 8.6$ $37.3\pm 22.9$ $0.9\pm 1.0$ $19.8\pm 0.1$
Credit Accuracy $98.2\pm 0.1$ $96.0\pm 0.5$ $44.3\pm 1.6$ $93.8\pm 0.2$ $53.8\pm 25.8$ $41.1\pm 2.8$ $13.9\pm 0.1$
Discrete Acc $98.5\pm 0.1$ $97.1\pm 0.5$ $75.2\pm 4.8$ $92.2\pm 0.3$ $76.7\pm 25.9$ $88.4\pm 5.7$ $10.6\pm 0.1$
Continuous Acc $98.0\pm 0.2$ $95.3\pm 0.6$ $26.6\pm 1.8$ $94.8\pm 0.2$ $51.9\pm 26.6$ $14.0\pm 3.4$ $15.7\pm 0.1$
Table 3: Performance comparison with state-of-the-art methods on the bank, adult, and credit datasets. (Best results are highlighted in bold.)
QA DPA IQA SA Black-box Query-free
PSNR $23.83$ $25.61$ $14.92$ $22.32$ $19.81$ $12.74$
SSIM $0.85$ $0.89$ $0.58$ $0.81$ $0.69$ $0.18$
Table 4: Performance comparison with data reconstruction methods used in split learning on the CIFAR10 dataset. Here, QA, DPA, IQA, and SA represent the results of UIFV in four different scenarios, while Black-box and Query-free are from comparative methods. (Best results are highlighted in bold.)
In our VFL data reconstruction attack experiments, we conducted a comprehensive comparison between our proposed UIFV method and state-of-the-art approaches on bank, adult, and credit datasets. Additionally, we compared UIFV against split learning methods on CIFAR-10. We conducted experiments using the UIFV method in four different scenarios, while for the comparison methods, we tested them only in their respective originally proposed specific scenarios. We utilized the evaluation metrics from Section 6.2 and set the epsilon value for continuous features to 0.2. Detailed results can be found in Tables 3 and 4.
The research findings indicate that our UIFV method significantly outperformed the comparison methods in terms of reconstruction accuracy across the four different scenarios. The QA scenario, with its relaxed scenario assumptions, excelled across datasets, surpassing 96% in overall, discrete, and continuous accuracy on bank and adult datasets. Its continuous accuracy reached 96.4%±0.2 on the bank dataset and 98.8%±0.1 on the adult dataset, outperforming other methods like GRN and GIA, which had much lower rates. Similarly, DPA, IQA, and SA scenarios showed robust performance, with most outcomes exceeding 60% accuracy. Compared to GRN and GIA, our methods demonstrated superior performance across all metrics. For instance, IQA and SA showed lower accuracies of 66.2%±1.4 and 90.1%±0.1 on the bank dataset, while GRN and GIA lagged further behind.
In the experimental section of Section 6, we conducted detailed reconstruction experiments on the image dataset, using two metrics, PSNR and SSIM, to measure the effectiveness of the reconstruction. However, while these metrics can assess our results from a technical standpoint, they are not sufficient to intuitively display the actual effects of the attack. Therefore, Figure 3 directly presents some specific reconstruction results, vividly demonstrating the actual effectiveness of our attack method.
We conducted reconstruction experiments on the CIFAR10 image dataset using two metrics, PSNR and SSIM, to measure the effectiveness of our UIFV method. In the DPA scenario, we achieved the highest scores with a PSNR of 25.61 and an SSIM of 0.89, demonstrating robustness in maintaining image quality. These values significantly surpassed those of comparative methods; the Black-box approach scored a PSNR of 19.81 and an SSIM of 0.69, while the Query-free method registered markedly lower scores of 12.74 and 0.18, respectively. Our UIFV method also showed commendable performance in other scenarios. To visually illustrate our attack outcomes, we present some actual reconstruction results in Figure 3.
@@
Figure 3: Our method is applied to the CIFAR10 dataset. The first line is the original image, the second and second lines are the private data that needs to be reconstructed during the VFL process, and the last four lines are the reconstruction effects under the four scenes.
@@ 6.3.1 Attack Effectiveness at Different Feature Splitting Ratios
The results, depicted in Figure 4, show that different methods performed variably across feature splitting ratios and datasets. Generally, QA, DPA, and SA methods yielded stable and effective results, with attack accuracy rates over 60%, indicating significant privacy risks. The IQA method was less stable but still outperformed random guessing.

### v2

@@ 1 Function StealthAttack($H^{\text{target}}$, $D^{\text{leak}}$):
3      $g$=TrainInverNetWithLeakedData($D^{\text{leak}}$)
4      $\hat{x}_{\text{target}}$ = Inverse($g$, $H^{\text{target}}$)
@@
8Function TrainInverNetWithLeakedData($D^{\text{leak}}$):
@@ 10      $g^{(0)}=$ Init()
13            Obtain leaked samples $x_{1}^{\text{target}},x_{2}^{\text{target}}...x_{m}^{\text{target}}$ from $D^{\text{leak}}$
14            Prepare $H_{i}^{\text{target}}$ data corresponding to $x_{i}^{\text{target}}$ , where $m\ll|H^{\text{target}}|$
15            $L(g)=\frac{1}{m}\sum_{i=1}^{m}\left\|g(H_{i}^{\text{target}})-x_{i}^{\text{target}}\right\|^{2}$
16            $\theta_{g}^{(n+1)}=\theta_{g}^{(n)}-\epsilon\frac{\partial L(g^{(n)})}{\partial\theta_{g}^{(n)}}$
@@ 17            $n+=1$
19      return $g^{(\text{NIters})}$
@@
In the previous three attack scenarios, the attacker can possess an i.i.d. auxiliary dataset and/or make queries to obtain the intermediate features but does not know about the target party’s private information. In this scenario, even if the attacker loses both the auxiliary data and the query capabilities, data reconstruction is still possible if the attacker $\mathbb{P}_{\text{adv}}$ may acquire a minimal amount of the target party’s private data $D^{\text{leak}}(D^{\text{leak}}\subset x^{\text{target}})$ and explicitly knows that these data have already been used in the VFL process. For instance, $\mathbb{P}_{\text{adv}}$ might collude with some internal employees of the target party (whose data are jointly maintained by both parties) to secretly acquire a small subset of data, which is considered possible as has been noted in some studies on analogous VFL scenarios, such as [32]. The knowledge of some private data used by the target party allows the attacker to start from the second stage of the UIFV framework to directly train InverNet $g$ with the objective function simplified to 9:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\arg\min_{\theta_{g}}\left\|g(H^{\text{target}}_{\text{sub}})-x^{\text{target}}_{\text{sub}}\right\|^{2}}$ (9)
In this equation, $x^{\text{target}}_{\text{sub}}$ signifies the secretly acquired data, while $H^{\text{target}}_{\text{sub}}$ represents the corresponding intermediate features of these secret data. Then, the trained $g$ will be used in the third stage to complete private data reconstruction. The complete attack algorithm in the Stealth Attack scenario is shown in Algorithm 4.
@@ 6.1.1 Datasets
In our experiments, we employed four public datasets: Bank marketing analysis [33] [^1] https://archive.ics.uci.edu/dataset/222/bank+marketing (Bank), Adult income [34] [^2] https://archive.ics.uci.edu/dataset/2/adult (Income), Default of credit card clients [35] [^3] https://archive.ics.uci.edu/dataset/350/default+of+credit+card+clients (Credit) and CIFAR10 [36] [^4] https://www.cs.toronto.edu/~kriz/cifar.html, to evaluate our methods. These datasets range from banking marketing analysis to image recognition, each with its unique features and challenges. For data preparation, continuous columns in tabular datasets were scaled, and discrete columns were one-hot encoded. We have summarized the evaluated datasets in Table 3.
Dataset Bank Income Credit CIFAR10
@@
The Bank dataset, derived from a Portuguese banking institution’s direct marketing campaigns, encompasses 41188 samples with 20 features, including 10 discrete features, aimed at determining the likelihood of clients subscribing to a term deposit. The features include age, job, marital status, and education. The Income dataset, also known as the ”Census Income” dataset, includes 32561 instances with 14 features, including 8 discrete features such as work class, and education, and it aims to predict if an individual’s income surpasses \$50,000 per annum using census data. The Credit dataset, sourced from Taiwan, comprises 30,000 instances with 23 features, including 9 discrete features such as credit amount, gender, education, and marital status, and aims to predict the probability of default payments in credit card clients using various data mining methods. The CIFAR10 dataset is a widely-used public dataset for computer vision research, containing 60,000 color images with a resolution of 32x32 pixels, divided into ten categories with 6,000 images in each category. Additionally, before training the models, we scaled the continuous columns in the tabular datasets to the range of [-1, 1], while the discrete columns were encoded using one-hot encoding. For image data, no preprocessing was performed. We used 80% of the tabular data to train the VFL model, with the remaining 20% serving as the target for data reconstruction attacks. For the CIFAR10 dataset, we used the training set consisting of 50,000 images to train the VFL model, and the test set with 10,000 images was used as the target for data reconstruction attacks.
@@ 6.1.2 Models
The InverNet for all bottom models is consistent with the architecture of the respective bottom model. For tabular data, the InverNet uses a three-layer fully connected neural network; for image data, the model employs two transposed convolutional layers, with a ReLU activation function between each layer. Model details are provided in table 4.
Bank Income Credit CIFAR10 Bottom MLP MLP MLP Conv2d(3$\to$32,kernel=3,padding=1) Model (input_dim,300,100,100) (input_dim,300,100,100) (input_dim,300,100,100) Conv2d(64$\to$64, kernel=3,padding=1) MaxPool2d(kernel=2,stride=2) InverNet MLP MLP MLP ConvTranspose2d(64$\to$64,kernel=3,padding=1) (100,100,300,input_dim) (100,100,300,input_dim) (100,100,300,input_dim) ConvTranspose2d(64$\to$3,kernel=3,padding=1)
Table 4: Model architectures for different datasets.
@@ 6.2 Evaluation Metrics
For image data, we adopt two widely recognized metrics: Peak Signal-to-Noise Ratio (PSNR) and Structural Similarity Index (SSIM) [37]. PSNR quantifies image errors by calculating the mean squared error between origin and attack images, with higher values indicating lower quality degradation. SSIM evaluates image quality based on structural information, brightness, and contrast, ranging from 0 to 1, with 1 indicating perfect similarity.
In evaluating tabular data, previous studies [6, 25] have used training loss or distance measures to assess reconstruction accuracy. However, these methods may not align with real attack scenarios, which focus on whether reconstructed categories match the actual ones. To address these issues, we adopted the metrics proposed in [23]. Considering the characteristics of tabular data, we separate the treatment of categorical and continuous features. For vector $x$ and its reconstruction vector $\hat{x}$, the accuracy metric is defined as follows:
$\text{accuracy}(x,\hat{x}):=\frac{1}{M+L}\bigg{(}\sum_{i=1}^{M}\mathbb{I}\{x^{D}_{i}=\hat{x}^{D}_{i}\}+\sum_{i=1}^{L}\mathbb{I}\{\hat{x}^{C}_{i}\in[x^{C}_{i}-\varepsilon,x^{C}_{i}+\varepsilon]\}\bigg{)},$
@@ 6.3 Performance Evaluation and Comparison
In our VFL data reconstruction attack experiments, we conducted a comprehensive comparison between the proposed UIFV method and the latest state-of-the-art methods, using datasets including Bank, Income, and Credit. We evaluated the UIFV method in four different scenarios and ensured that the compared methods utilized the same architecture and consistent experimental settings as UIFV. For the GIA method [7] and the Ginver method [22], we adopted black-box attack versions where the attacker does not know the specific structure and parameters of the model. Additionally, we included a random guessing baseline method to evaluate the inherent performance of random predictions. During the evaluation, we applied the metrics defined in Section 6.2, setting the $\varepsilon$ value for continuous features to 0.2 and the batch size to 64. Detailed results can be found in Table 5.
Method Bank Income Credit DGL[18] $13.63\pm 1.00$ $33.20\pm 0.74$ $16.88\pm 1.80$ SQR[19] $27.53\pm 0.36$ $22.05\pm 0.50$ $15.50\pm 0.82$ CPA[20] $29.00\pm 2.83$ $33.33\pm 2.49$ $17.27\pm 2.26$ LOKI[21] $14.72\pm 0.74$ $0.34\pm 0.22$ $00.00\pm 0$ 0.00 GRN[6] $30.20\pm 6.68$ $41.12\pm 8.52$ $53.82\pm 25.84$ GIA[7] $55.93\pm 1.81$ $18.87\pm 8.27$ $41.07\pm 2.78$ Ginver[22] $78.23\pm 2.23$ $80.91\pm 2.07$ $69.44\pm 1.92$ Random $21.02\pm 0.04$ $11.14\pm 0.07$ $13.88\pm 0.06$ UIFV-QA $97.96\pm 0.11$ $98.49\pm 0.04$ $98.19\pm 0.13$ UIFV-DPA $95.74\pm 0.38$ $80.80\pm 0.91$ $96.00\pm 0.46$ UIFV-IQA $66.17\pm 1.43$ $94.81\pm 0.54$ $44.25\pm 1.61$ UIFV-SA $90.07\pm 0.11$ $72.79\pm 3.40$ $93.83\pm 0.20$
[…]
Due to differences in attack assumptions among the methods (as detailed in Table 1), it is challenging to directly compare our method with others. However, overall, the UIFV method achieved relatively high attack success rates under the weakest attack assumptions. When conducting a lateral comparison within specific scenarios, only GIA and Ginver overlap with UIFV in terms of attack scenarios: the experimental setting of GIA aligns with the UIFV-DPA scenario, but its performance across the three datasets is significantly lower than that of UIFV. Ginver’s experimental scenario was more similar to UIFV-IQA, and their performances were comparable, with each method excelling in different aspects. However, Ginver was less flexible and applicable than UIFV.
We also observed that gradient-based attack methods, such as DGL and SQR, performed poorly across the three datasets, primarily due to the batch size being set to 64, which significantly impacted their attack performance. Similarly, model information-based attack methods, such as GRN and GIA, also exhibited poor performance on tabular data. This is because tabular data typically features high dimensionality and low correlation, making the optimization problem non-convex and complex, which often leads to local optima and reduces the likelihood of successful optimization.
6.3.1 Performance on Tabular Datasets
Dataset Evaluation UIFV-QA UIFV-DPA UIFV-IQA UIFV-SA
Bank Accuracy $98.0\pm 0.1$ $95.7\pm 0.4$ $66.2\pm 1.4$ $90.1\pm 0.1$
Discrete Acc $99.5\pm 0.0$ $99.0\pm 0.1$ $87.8\pm 2.1$ $94.9\pm 0.3$
Continuous Acc $96.4\pm 0.2$ $92.4\pm 0.8$ $44.6\pm 1.2$ $85.2\pm 0.3$
Income Accuracy $98.5\pm 0.0$ $80.8\pm 0.9$ $94.8\pm 0.5$ $72.8\pm 3.4$
Discrete Acc $98.4\pm 0.0$ $81.1\pm 0.9$ $98.1\pm 0.6$ $71.4\pm 2.4$
Continuous Acc $98.8\pm 0.1$ $79.8\pm 1.2$ $81.7\pm 0.7$ $78.3\pm 8.6$
Credit Accuracy $98.2\pm 0.1$ $96.0\pm 0.5$ $44.3\pm 1.6$ $93.8\pm 0.2$
Discrete Acc $98.5\pm 0.1$ $97.1\pm 0.5$ $75.2\pm 4.8$ $92.2\pm 0.3$
Continuous Acc $98.0\pm 0.2$ $95.3\pm 0.6$ $26.6\pm 1.8$ $94.8\pm 0.2$
Table 6: Reconstruction Performance of UIFV on Discrete and Continuous Features Across Four Scenarios.
To further analyze the performance of UIFV in tabular data reconstruction, we divided the data from three datasets into two categories: discrete data and continuous data. As shown in Table 6, UIFV achieves significantly higher reconstruction accuracy on discrete data (encoded with one-hot) compared to continuous data across all four scenarios. For discrete data, UIFV consistently achieves over 70% reconstruction accuracy in all scenarios. For continuous data, except in the IQA scenario, UIFV also demonstrates over a 70% probability of reconstructing values close to the original. This highlights a significant privacy threat to VFL systems.
6.3.2 Performance on Image Datasets
UIFV-QA UIFV-DPA UIFV-IQA UIFV-SA
PSNR $23.83$ $25.61$ $14.92$ $22.32$
SSIM $0.85$ $0.89$ $0.58$ $0.81$
Table 7: Reconstruction Performance of UIFV on the CIFAR10 dataset Across Four Scenarios.
To further demonstrate the generalizability of the UIFV method, we conducted reconstruction experiments on the CIFAR-10 image dataset, using PSNR and SSIM as metrics to evaluate the effectiveness of UIFV. The experimental results are shown in Table 7. The results indicate that UIFV performs slightly better in the DPA scenario, which we attribute to the high spatial correlation in image data. This correlation benefits InverNet during shadow model training by facilitating the capture of relationships between intermediate features and the original image. Overall, UIFV exhibited strong attack performance across all four scenarios, posing a significant threat to VFL security. To provide a more intuitive understanding of our attack results, we present some actual reconstructed images in Figure 6.
@@ 6.3.2 Performance on Image Datasets
Figure 6: Our method is applied to the CIFAR10 dataset. The first line is the original image, the second and second lines are the private data that needs to be reconstructed during the VFL process, and the last four lines are the reconstruction effects under the four scenarios.
@@ 6.3.3 Attack Effectiveness at Different Feature Splitting Ratios
The results, depicted in Fig. 7, show that different methods performed variably across feature splitting ratios and datasets. Generally, QA, DPA, and SA methods yielded stable and effective results, with attack accuracy rates over 60%, indicating significant privacy risks. The IQA method was less stable but still outperformed random guessing.
6.4 Defense evaluation
Bank Income
Ratio AUC QA DPA IQA SA Ratio AUC QA DPA IQA SA
$1$ $0.859$ $86.38$ $88.29$ $18.96$ $18.96$ $1$ $0.888$ $87.78$ $84.92$ $73.92$ $73.87$
$0.5$ $0.938$ $97.46$ $95.40$ $52.80$ $18.96$ $0.5$ $0.877$ $92.27$ $87.59$ $84.15$ $83.07$
$0.1$ $0.938$ $97.93$ $95.50$ $64.01$ $78.16$ $0.1$ $0.876$ $93.51$ $89.29$ $82.91$ $81.45$
$0.01$ $0.938$ $98.31$ $96.02$ $69.68$ $89.48$ $0.01$ $0.870$ $95.44$ $88.56$ $82.94$ $87.50$
$0.001$ $0.939$ $98.22$ $96.04$ $66.40$ $90.49$ $0.001$ $0.867$ $95.86$ $89.15$ $61.72$ $85.48$
Credit CIFAR10
Ratio AUC QA DPA IQA SA Ratio ACC QA DPA IQA SA
$1$ $0.770$ $96.75$ $92.87$ $28.92$ $88.46$ $1$ $61.99$ $0.70$ $0.85$ $0.34$ $0.11$
$0.5$ $0.760$ $96.52$ $96.52$ $26.04$ $90.45$ $0.5$ $67.26$ $0.14$ $0.81$ $0.14$ $0.0$
$0.1$ $0.774$ $97.95$ $96.00$ $39.92$ $93.13$ $0.1$ $72.99$ $0.72$ $0.87$ $0.44$ $0.68$
$0.01$ $0.771$ $97.59$ $96.25$ $41.50$ $93.26$ $0.01$ $74.84$ $0.80$ $0.88$ $0.53$ $0.74$
$0.001$ $0.765$ $98.36$ $95.09$ $41.05$ $93.29$ $0.001$ $74.25$ $0.86$ $0.89$ $0.57$ $0.81$
Bank Income
Ratio AUC QA DPA IQA SA Ratio AUC QA DPA IQA SA
$1$ $0.937$ $18.96$ $89.19$ $49.02$ $18.96$ $1$ $0.885$ $70.87$ $69.93$ $56.21$ $3.07$
$0.5$ $0.939$ $98.52$ $94.51$ $64.81$ $18.96$ $0.5$ $0.892$ $91.20$ $71.88$ $90.92$ $3.07$
$0.1$ $0.940$ $98.22$ $95.48$ $67.16$ $90.67$ $0.1$ $0.870$ $96.37$ $90.68$ $73.40$ $87.19$
$0.01$ $0.940$ $97.85$ $95.59$ $63.60$ $88.81$ $0.01$ $0.886$ $91.14$ $82.20$ $72.69$ $83.50$
$0.001$ $0.939$ $98.26$ $96.20$ $60.42$ $91.19$ $0.001$ $0.868$ $94.90$ $89.17$ $78.52$ $86.16$
Credit CIFAR10
Ratio AUC QA DPA IQA SA Ratio ACC QA DPA IQA SA
$1$ $0.769$ $98.13$ $87.77$ $62.98$ $3.46$ $1$ $64.97$ $0.75$ $0.79$ $0.59$ $0.0$
$0.5$ $0.775$ $98.35$ $96.08$ $57.51$ $94.34$ $0.5$ $66.81$ $0.88$ $0.79$ $0.62$ $0.07$
$0.1$ $0.769$ $98.04$ $96.20$ $33.98$ $91.86$ $0.1$ $74.41$ $0.86$ $0.83$ $0.59$ $0.82$
$0.01$ $0.772$ $97.89$ $96.24$ $37.66$ $91.96$ $0.01$ $74.13$ $0.81$ $0.87$ $0.56$ $0.76$
$0.001$ $0.768$ $97.54$ $95.35$ $33.91$ $92.52$ $0.001$ $74.38$ $0.80$ $0.90$ $0.57$ $0.90$
Table 9: Experimental Results of Gaussian noise Defense.
Although this study primarily focuses on attack strategies, we have also examined several defensive measures. We first considered two common defensive measures: Differential Privacy (DP) and noise addition. Differential Privacy technology defends against data reconstruction attacks by adding noise to gradients during the training process, thereby protecting individual data privacy while maintaining the overall effectiveness of the model. The method of adding noise to model outputs hinders attackers from obtaining precise information, thus protecting the data from malicious use, although this may affect the accuracy of the model.
To test the effectiveness of these defensive methods, we set the noise ratio to 1, 0.5, 0.1, 0.01, and 0.001, respectively, and conducted experiments across four different scenarios in four datasets. The experimental results for Differential Privacy are shown in Table 8, and those for noise addition are shown in Table 9. These two defensive methods indeed reduce the effectiveness of data reconstruction attacks to some extent. However, their impact on attack effectiveness is relatively limited. In scenarios such as QA, DPA, and IQA, attacks can still maintain a certain success rate even with a high noise ratio. In the SA scenario, the impact of these two defensive methods is more noticeable. As the noise ratio increases, there is a downward trend in attack accuracy in the SA scenario.
To further investigate the effectiveness of defense methods, we implemented two approaches designed to address privacy leakage in Federated Learning (FL). The first method is PA-iMFL[38], a privacy amplification approach targeting data reconstruction attacks in advanced multi-layer federated learning. By combining local differential privacy, privacy-enhanced subsampling, and gradient sign resetting, PA-iMFL achieves bidirectional gradient compression, which not only improves communication efficiency but also strengthens privacy protection. The second method is VFLDefender[39], which protects privacy by disrupting the correlation between gradients and training samples during model updates, thereby reducing attackers’ ability to reconstruct labels or features.
Dataset AUC/ACC QA DPA IQA SA
Bank 0.925 98.65 92.98 80.17 94.08
Income 0.893 89.55 77.24 73.82 58.32
Credit 0.767 98.50 92.72 70.73 96.93
CIFAR10 68.64 0.640 0.896 0.220 0.218
Table 10: Experimental Results of PA-iMFL Defense. For each dataset, the first column shows the results of the VFL task, and the last four columns indicate the effectiveness of our attack method under four different scenarios. For CIFAR10, SSIM is the Evaluation Metric.
Dataset AUC/ACC QA DPA IQA SA
Bank 0.865 98.67 93.04 80.60 89.11
Income 0.891 78.53 75.86 77.27 83.72
Credit 0.767 98.63 94.24 65.80 97.00
CIFAR10 67.01 0.226 0.865 0.148 0.203
Table 11: Experimental Results of VFLDefender Defense.
We evaluated the defensive effects of PA-iMFL and VFLDefender on four datasets. The experimental results for PA-iMFL are shown in Table 10, while those for VFLDefender are shown in Table 11. The results demonstrate that both defense methods have a minimal impact on the primary VFL training tasks and effectively reduce the efficacy of data reconstruction attacks to some extent, particularly showing significant defensive effects on image datasets. However, neither method provided effective defense in the DPA scenario. This is primarily because, in the DPA scenario, the attack does not require any query requests to the target party, rendering communication-focused defense measures partially ineffective.